OpenStudy (anonymous):

find all the real zeros of the polynomial function. determine the multiplicity of each zero. f(x)=2x^4-2x^2-40

OpenStudy (campbell_st):

this is an equation reducible to a quadratic let u = x^2 or just factorise \[(x^2 - 5)(2x^2 + 8) = 0\] then \[x^2 = 5... and .... x^2 = -8\] I'll leave you to identify the roots from here... 2 are irrational and 2 are complex